OpenStudy (anonymous):

What is the value of the x determinant (Dx) in the system? 3x – 7y = 11 2x + 5y = –4

OpenStudy (anonymous):

multiply equation one by 2 and equation two by 3 and subtract equation two from one . solve for y, substitute that value of y to obtain values of x

OpenStudy (anonymous):

\[ax + by = k_1\] \[cx+ dy = k_2\] \[determinant = ad -bc\]

OpenStudy (anonymous):

D(x) is being asked for

OpenStudy (anonymous):

so it makes sense to obtain eigen values : )

OpenStudy (anonymous):

solve for either a row or column vector

OpenStudy (anonymous):

det x = 27 / 29. Test it and see. also you will find det y = -34 / 29.

OpenStudy (anonymous):

sorry, misread another question. i think im getting tired